OpenStudy (anonymous):

Danely has only nickels and quarters in her piggy bank. Their combined value is $9.15. Their combined weight is one pound. Ninety nickels weigh one pound. Eighty quarters weight one pound. How many nickels does Danely have in her piggy bank?

OpenStudy (anonymous):

Well let's set up 2 equations: We know that a quarter is worth .25, and a nickel is worth .05, so our first equation is going to look like .25x+.05y=9.15 The second equation is a little tricky, but if you have 80 quarters, you have 1 pound right? so 80x=1 or x=1/80 90 nickels is also a pound, so y=1/90. we want the total to be 1 pound, so our equation could be the weight of a single coin, times how many coins you have to equal 1 pound, so (1/80)x+(1/90)y=1 .25x+.05y=9.15 All you need to do is solve for one variable, plug it in to the other equation, and solve for the other variable. hope this helps!